### v1

# Goldfish: Monolingual Language Models for 350 Languages

Abstract
For many low-resource languages, the only available language models are large multilingual models trained on many languages simultaneously. However, using FLORES perplexity as a metric, we find that these models perform worse than bigrams for many languages (e.g. 24% of languages in XGLM 4.5B; 43% in BLOOM 7.1B). To facilitate research that focuses on low-resource languages, we pre-train and release Goldfish, a suite of monolingual autoregressive Transformer language models up to 125M parameters for 350 languages. The Goldfish reach lower FLORES perplexities than BLOOM, XGLM, and MaLA-500 on 98 of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller. However, the Goldfish significantly underperform larger multilingual models on reasoning benchmarks, suggesting that for low-resource languages, multilinguality primarily improves general reasoning abilities rather than basic text generation. We release models trained on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) of text data where available. The Goldfish models are available as baselines, fine-tuning sources, or augmentations to existing models in low-resource NLP research, and they are further useful for crosslinguistic studies requiring maximally comparable models across languages.
Goldfish: Monolingual Language Models for 350 Languages
[…]
1 Introduction
Language modeling research in low-resource languages often relies on large multilingual models trained on many languages simultaneously (Conneau et al., 2020; Adelani et al., 2021b; Ebrahimi et al., 2022; Lin et al., 2022; Hangya et al., 2022; Imani et al., 2023). For many low-resource languages, a dedicated model optimized for that language does not exist. This lack of dedicated models hinders comparability of results across models and languages (Bandarkar et al., 2024), and it contributes to model under-performance in low-resource languages (Wu and Dredze, 2020; Blasi et al., 2022). These barriers to research in low-resource languages are likely to exacerbate existing inequities across language communities in NLP research (Bender, 2011; Joshi et al., 2020).
To address this lack of available models, we introduce Goldfish, a suite of over 1000 monolingual language models for 350 diverse languages. [^1] The name refers to shared qualities between our models and goldfish (Carassius auratus); they are small, there are many of them, and they are known for their poor memories (perhaps inaccurately; Carey, 2024). The models reach lower perplexities than XGLM (Lin et al., 2022), BLOOM 7.1B (Scao et al., 2022), and MaLA-500 (Lin et al., 2024) on 98 out of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller. The Goldfish also outperform simple bigram models, which are surprisingly competitive with larger models for low-resource languages (e.g. lower perplexities than BLOOM 7.1B on 43% of its languages; §4). However, despite better perplexities, the Goldfish underperform larger multilingual models on reasoning benchmarks, suggesting that multilingual pre-training may benefit abstract reasoning capabilities over more basic grammatical text generation (§5).
Data size Model output 5MB ‘‘Goldfish are a few years of the most of the most of the most...’’ 10MB ‘‘Goldfish are a great way to the best way to the best way...’’ 100MB ‘‘Goldfish are a great way to get your fish in the wild.’’ 1GB ‘‘Goldfish are a species of fish that are found in the sea.’’
Figure 1: Left: Map of the 350 languages for which Goldfish models are available, using coordinates from Glottolog (Hammarström et al., 2023). Right: Sample model outputs completing the prompt ‘‘Goldfish are’’ for the eng_latn (English) model for each dataset size, using sampling temperature zero. Grammatical text generation begins to emerge in the 100MB-dataset model (available for 166 languages), but the lower-resource models still achieve better perplexities than previous models for many low-resource languages (§4).
Finally, to enable comparisons across languages, we release monolingual models trained on comparable dataset sizes for all languages: 5MB, 10MB, 100MB, and 1GB when available, after accounting for the fact that languages require different numbers of UTF-8 bytes to encode comparable content (Arnett et al., 2024). These Goldfish serve as baselines, allowing results in diverse languages to be situated relative to comparable models. They can also be used as source models for fine-tuning or to enhance larger multilingual models in areas where those models fall short (§4). Models and code are available at https://huggingface.co/goldfish-models.
2 Related Work
Low resource language modeling often leverages multilingual pre-training, where a model is trained on multiple languages simultaneously (Pires et al., 2019; Conneau et al., 2020). Indeed, this can improve low-resource performance, particularly when models have sufficient capacity and the multilingual data is from related or typologically similar languages (Kakwani et al., 2020; Ogueji et al., 2021; Chang et al., 2023). However, monolingual models have still been shown to achieve better performance than multilingual models for many languages (e.g. Martin et al., 2020; Pyysalo et al., 2021; Gutiérrez-Fandiño et al., 2021; Luukkonen et al., 2023). Thus, it appears that existing multilingual language models are still limited by model capacity or limited data in low-resource languages (Conneau et al., 2020; Chang et al., 2023).
Notably, the training datasets for massively multilingual models are often heavily skewed towards high-resource languages. For example, XGLM 4.5B is trained on over 7000$\times$ more Norwegian (71GB; 5.4M native speakers) than Quechua (0.01GB; 7.3M native speakers; Lin et al., 2022; Ethnologue, 2024). In a more extreme case, BLOOM is trained on only 0.07MB of Akan (8.1M native speakers) out of 1.61TB total (4e-6% of the pre-training dataset; Scao et al., 2022). These extremely small quantities of low-resource language data often do not leverage recent efforts to compile text data in low-resource languages (Costa-jussà et al., 2022; Imani et al., 2023; Kudugunta et al., 2023), and the data imbalances are likely to severely hinder performance in low-resource languages. Indeed, we find that these models have worse perplexities than simple bigram models for many languages (§4). Unfortunately, comparable monolingual language models across many diverse languages have yet to be studied or released.
3 Models and Datasets
We introduce the Goldfish models, a suite of 1154 monolingual Transformer language models pre-trained for 350 languages. The largest model for each language is 125M parameters. We train models on 5MB, 10MB, 100MB, and 1GB of text when available after byte premium scaling (Arnett et al., 2024). Figure 1 shows a geographic map of the 350 languages, with coordinates from Glottolog (Hammarström et al., 2023), along with sample outputs from the English model for each dataset size.
Goldfish data size # Langs Goldfish Bigrams XGLM 4.5B MaLA-500 10B
1000MB 73 76.9 112.3 78.6 84.7
100MB 22 102.7 132.6 143.9 121.7
10MB, 5MB 5 130.5 148.3 183.1 135.0
Table 1: Mean FLORES log-perplexity ($\downarrow$) for the 100 languages in XGLM 4.5B, MaLA-500, and FLORES, separated by maximum Goldfish dataset size. The Goldfish languages are a strict superset of these languages.
3.1 Training Datasets
We merge the massively multilingual text datasets compiled in Chang et al. (2023), Glot500 (Imani et al., 2023), and MADLAD-400 (Kudugunta et al., 2023) per language. To facilitate fair evaluations, we hold out FLORES-200 and AmericasNLI from all datasets (Costa-jussà et al., 2022; Ebrahimi et al., 2022). We deduplicate repeated sequences of 100 UTF-8 bytes and drop languages with only Bible data. Full dataset details are in §A.1.
To sample pre-training datasets of the desired sizes in a language $L$, we first use the Byte Premium Tool (Arnett et al., 2024) to estimate the byte premium for $L$, the number of UTF-8 bytes required to encode comparable text in $L$ relative to eng_latn (English). For example, khm_khmr (Khmer) has byte premium 3.91, meaning that it uses approximately 3.91$\times$ as many UTF-8 bytes as English to encode content-matched text. We divide each dataset size by the estimated byte premium for the corresponding language, thus measuring all datasets in units of “equivalent” English text bytes. We sample datasets to train monolingual language models on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) when available after byte premium scaling. [^2] The languages with 5MB-dataset models are a subset of the languages with 10MB-dataset models, and similarly for the 100MB and 1GB dataset sizes. These are equivalent to roughly 1M, 2M, 20M, and 200M tokens of English text respectively; including 10 epochs of repetition, the 1GB-dataset models are trained on the equivalent of roughly 2B English tokens. When a 1GB dataset is not available for a language after byte premium scaling, we include a full model (267 languages) trained on the entire dataset in that language, for use cases that seek to maximize performance in a specific low-resource language.
3.2 Architectures and Pre-Training
For each language and each dataset size, we pre-train an autoregressive GPT-2 Transformer language model from scratch (Radford et al., 2019). For the 1GB, 100MB, and full dataset sizes, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018), which has a similar parameter count to BERT-base and RoBERTa (Devlin et al., 2019; Liu et al., 2019). Because larger models do not appear to outperform smaller models for very small datasets (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes. Full hyperparameters are reported in §A.2.
We tokenize each dataset using using a monolingual SentencePiece tokenizer (Kudo and Richardson, 2018) trained on that dataset size, limiting tokenizer training text to 100MB after byte premium scaling. Following Liu et al. (2019), we use vocabulary size 50K and a maximum sequence length of 512 tokens for all models. We train each language model on 10 epochs of its corresponding dataset. [^3] Multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but we find that more than 10 epochs of training leads to overfitting for extremely small datasets (e.g. 5MB). Pre-training details, compute costs, and all available models are reported in §A.2.
Bigrams XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
Bigrams 24 / 102 0 / 30 20 / 46 11 / 175
Goldfish (ours) 202 / 202 60 / 102 2 / 30 32 / 46 111 / 175
Table 2: FLORES perplexity win rates for each row vs. column model. For example, Goldfish reach lower log-perplexities than MaLA-500 for 111/175 (63%) of FLORES languages in both Goldfish and MaLA-500.
# Langs Chance Goldfish XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
Belebele 121 25.0 28.2 30.1 30.6 30.2 30.6
XCOPA 11 50.0 54.9 57.9 60.6 56.9 55.6
XStoryCloze 10 50.0 52.5 57.1 59.9 58.2 55.7
Table 3: Reasoning benchmark accuracies averaged over non-English languages. Despite better perplexities, the Goldfish perform significantly worse than larger multilingual models on reasoning.
4 FLORES Log-Perplexity Evaluations
We first evaluate our models on FLORES-200 log-perplexity (Costa-jussà et al., 2022) (equivalently, negative log-likelihood; Lin et al., 2024). To avoid tokenization confounds from computing log-perplexity per token, we compute log-perplexity per FLORES sequence. Regardless of its tokenization, a language model $\mathcal{M}$ assigns some probability $P_{\mathcal{M}}(s)$ to each sequence $s$ in FLORES. In most cases, $s$ is a single sentence. For fair comparison with multilingual models that need to determine the input language during the early parts of a sequence, we compute log-perplexity of the second half $s_{1}$ of each sequence given the first half $s_{0}$. We then compute the mean over sequences:
$\textrm{LogPPL}_{\mathcal{M}}=\textrm{mean}_{s}\Big{(}-\textrm{log}(P_{\mathcal{M}}(s_{1}|s_{0}))\Big{)}$ (1)
A lower log-perplexity indicates better performance, where $\mathcal{M}$ assigns higher probabilities to ground truth text (FLORES sequences). While imperfect, perplexity does not require annotated text data, it is predictive of performance on a variety of downstream tasks (Xia et al., 2023), and it has been used to measure language model quality in previous work (Kaplan et al., 2020; Hoffmann et al., 2022; Lin et al., 2024).
We compare the Goldfish models with XGLM 4.5B (Lin et al., 2022; 134 languages), XGLM 7.5B (30 languages), BLOOM 7.1B (Scao et al., 2022; 46 languages), and MaLA-500 10B (Lin et al., 2024; 534 languages). We also compare to simple bigram models trained on the Goldfish datasets. [^4] Bigram and perplexity implementation details in §A.3. In all cases, we use the Goldfish model trained on the maximum amount of data in each language (maximum 1GB).
FLORES log-perplexity results.
The Goldfish reach lower log-perplexities than all four comparison models on 98 of the 204 FLORES languages. Average log-perplexities for the 100 FLORES languages included in both XGLM 4.5B and MaLA-500 are reported in Table 1 (excluding XGLM 7.5B and BLOOM 7.1B there because they are trained on far fewer languages). On average, the Goldfish reach 13% lower log-perplexities than XGLM 4.5B, and 11% lower than MaLA-500 10B.
To ensure that these results are not driven by a small subset of specific languages, in Table 2 we report the pairwise “win” rates for Goldfish and bigrams vs. all four comparison models, for the set of FLORES languages shared between each pair. The Goldfish models have a perplexity win rate above 50% against all comparison models except XGLM 7.5B, which considers only 30 fairly high-resource languages (Lin et al., 2022). Notably, the bigram models also reach lower perplexities than large multilingual models for a nontrivial number of languages: 24% of languages in XGLM 4.5B and 43% of languages in BLOOM 7.1B. Still, the bigrams have worse perplexities than Goldfish for all languages. Log-perplexities for individual languages and models are reported in Table LABEL:tab:ppl-table.
5 Multilingual Reasoning Benchmarks
Because FLORES perplexities are not necessarily reflective of complex capabilities in language models, we also evaluate Goldfish, XGLM, BLOOM, and MaLA-500 (as in §4) on non-English Belebele (121 languages, reading comprehension; Bandarkar et al., 2024), XCOPA (11 languages, commonsense; Ponti et al., 2020), and XStoryCloze (10 languages, story commonsense; Lin et al., 2022). All models are evaluated zero shot with no fine-tuning. Evaluation task details are in §A.4.
Results for all three reasoning tasks are reported in Table 3. Although all models perform quite poorly (close to chance accuracy), the Goldfish perform substantially worse than the multilingual models. [^5] It is unlikely that this effect is due to model size alone; the Goldfish models (125M parameters) have easily enough capacity for their maximum of 1GB of text data. This indicates that the combination of larger datasets and model sizes in multilingual pre-training can allow language models to develop reasoning capabilities in specific languages, even when perplexities in those languages remain high. For example, XGLM 7.5B has worse perplexities than Goldfish for 82 Belebele languages (in fact, worse than bigrams for 77 languages), but it outperforms Goldfish on Belebele (reading comprehension) for 56 of those languages. This is in stark contrast with monolingual language models, which generally must reach low perplexities and acquire basic grammatical capabilities before developing reasoning abilities (Liu et al., 2021; Choshen et al., 2022; Xia et al., 2023; Chang et al., 2024). Intuitively, it may be that abstract reasoning patterns are often more language-agnostic than grammatical text generation, and thus multilingual pre-training primarily benefits the former.
6 Conclusion
We pre-train and release Goldfish, a suite of over 1000 monolingual language models for 350 languages. The Goldfish achieve perplexities that are competitive with, and on average lower than, state-of-the-art multilingual language models across languages. However, they underperform large multilingual models on reasoning tasks; in low-resource languages, it appears that multilingual pre-training facilitates nontrivial reasoning capabilities despite extremely poor perplexities. We publicly release all Goldfish models to be used as comparable baselines, fine-tuning sources, or augmentations to larger models (e.g. cross-lingual experts; Blevins et al., 2024) in future low-resource NLP research.
Limitations
Comparability and availability.
In order to include as many low-resource languages as possible, the Goldfish models are trained on corpora compiled from a wide variety sources (§A.1). Still, 5MB of text (roughly 1M tokens) is not publicly available for many of the world’s languages. Even where text is available, corpora for different languages vary significantly both in cleanliness and domain coverage (e.g. news vs. social media vs. books). Thus, while we release models trained on comparable quantities of text in different languages (including accounting for byte premiums; Arnett et al., 2024; §3.1), the models are not perfectly comparable across languages. In fact, it is likely that such perfect comparability is impossible given the diversity of the world’s languages, cultures, and language use. Even directly translated datasets are not perfectly comparable across languages (Jill Levine and Lateef-Jan, 2018). Thus, the Goldfish models aim to maximize model and dataset comparability across languages while still covering a wide variety of languages.
Monolinguality.
By design, all of the Goldfish models are monolingual. For low-resource languages, training on closely related languages would likely improve performance (Conneau et al., 2020; Chang et al., 2023). However, adding multilingual data introduces concerns such as the choice of added languages (some languages have more closely related languages in our dataset than others), quantities of added data, and model capacity limitations. To maximize comparability across languages and to allow the models to serve as clearly-defined baselines, we train all Goldfish models monolingually. Of course, language-annotated text datasets inevitably contain mislabeled text, particularly for similar languages (Caswell et al., 2020; Blevins and Zettlemoyer, 2022; Kreutzer et al., 2022). Thus, we cannot guarantee that our models are entirely free from cross-language contamination, although they are monolingual to the best ability of current language identification models.
Model and dataset sizes.
Because the Goldfish are focused on low-resource languages, we restrict all models to 1GB of training text (after byte premium scaling; Arnett et al., 2024). For the majority of the world’s languages, 1GB is sufficient to include all publicly available text data in the language. At these small dataset sizes, larger models do not appear to provide significant benefit over smaller models (Kaplan et al., 2020; Hoffmann et al., 2022; Chang et al., 2023). Thus, the largest Goldfish model that we train for each language has 125M parameters and is trained on a maximum of 1GB of text. This is the same model size as GPT-1 (Radford et al., 2018) or BERT (Devlin et al., 2019), and the 1GB dataset size is approximately 20% of the dataset size of GPT-1 (Radford et al., 2018).
Downstream tasks.
We evaluate the Goldfish models on FLORES log-perplexity (§4) and three reasoning benchmarks (§5). These are some of the only evaluations that can be used for autoregressive language models in many languages, but they have significant limitations. Perplexity is not necessarily predictive of grammatical text generation (Hu et al., 2020) or complex reasoning capabilities (Levy et al., 2024), but it still provides reasonable signal for model performance (Xia et al., 2023) and it is often used to roughly quantify language model quality (Kaplan et al., 2020; Hoffmann et al., 2022). On the other hand, reasoning benchmarks require annotated datasets and thus often cover fewer languages. One notable exception is Belebele (121 non-English languages; Bandarkar et al., 2024), but even large state-of-the-art models perform quite poorly on Belebele without tuning or few-shot prompting (§5). Thus, our evaluations of model reasoning are not entirely conclusive; we may primarily be measuring heuristics that allow the models to perform only somewhat above chance (arguably, this might still be considered a basic form of “reasoning”). We hope that tractable evaluation datasets with broad language coverage will become increasingly available in the future.
Risks and dataset licensing.
Trained on a maximum of 1GB of text each, the Goldfish models have very limited capabilities relative to modern language models in high-resource languages. The Goldfish are trained on publicly-released corpora used in previous NLP research (§A.1), but we cannot guarantee that the data is free from offensive content or personally identifying information. We do not redistribute the data itself. Furthermore, our models are small, which reduces the likelihood that they will regurgitate memorized text (Carlini et al., 2023). As far as we are aware, we do not include any datasets that prohibit use for language model training. We report all included datasets in §A.1. We will remove models for affected languages if contacted by dataset owners.
Acknowledgments
We would like to thank the UCSD Language and Cognition Lab for valuable discussion. Some models were trained on hardware provided by the NVIDIA Corporation as part of an NVIDIA Academic Hardware Grant. Some models were also trained on the UCSD Social Sciences Research and Development Environment (SSRDE). Zhuowen Tu is supported by NSF IIS-2127544. Tyler Chang is partially supported by the UCSD HDSI graduate fellowship.
Appendix A Appendix
A.1 Training Dataset Details
Data sources.
As described in §3.1, we merge the text datasets compiled in Chang et al. (2023), Glot500 (Imani et al., 2023), and MADLAD-400 (clean split; Kudugunta et al., 2023). These datasets include popular multilingual corpora such as OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), and others. Together, these datasets take advantage of both automatically crawled datasets with automated language identification and targeted datasets manually annotated for specific low-resource languages. All included datasets are publicly available; see Limitations for licensing concerns. Comprehensively, the Goldfish dataset includes:
  • Chang et al. (2023):
    OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), Leipzig Corpora Collection (Goldhahn et al., 2012), eBible translations (eBible, 2023), Tatoeba (Tiedemann, 2012, 2020), AfriBERTa (Ogueji et al., 2021), NusaX (Winata et al., 2023), AmericasNLP (Mager et al., 2021), Nunavut Hansard Inuktitut–English Parallel Corpus (Joanis et al., 2020), Cherokee-English ChrEn dataset (Zhang et al., 2020), Cherokee Corpus (Cherokee Corpus, 2023), Cree Corpus (Teodorescu et al., 2022), Languages of Russia (Zaydelman et al., 2016), Evenki Life newspaper (Zueva et al., 2020), transcribed Fula Speech Corpora (Cawoylel, 2023), IsiXhosa (Podile and Eiselen, 2016), Ewe Language Corpus (Gbedevi Akouyo et al., 2021), Makerere Luganda Corpora (Mukiibi et al., 2022), CMU Haitian Creole dataset (CMU, 2010), Tigrinya Language Modeling Dataset (Gaim et al., 2021), and Ulukau (Ulukau, 2023).
  • Glot500 (Imani et al., 2023):
    AI4Bharat (AI4Bharat, 2023), AI FOR THAI LotusCorpus (AI FOR THAI, 2023), Arabic Dialects Dataset (El-Haj et al., 2018), AfriBERTa (Ogueji et al., 2021), AfroMAFT (Adelani et al., 2022; Xue et al., 2021), Anuvaad (Anuvaad, 2023), AraBench (Sajjad et al., 2020), Autshumato (Autshumato, 2023) Bloom Library (Leong et al., 2022), CC100 (Conneau et al., 2020), CCNet (Wenzek et al., 2020), CMU Haitian Creole (CMU, 2010), SADiLaR NCHLT corpus (SADiLaR, 2023b), Clarin (Clarin, 2023), DART (Alsarsour et al., 2018), Earthlings (Dunn, 2020), FFR Dataset (FFR Dataset, 2023), GiossaMedia (Góngora et al., 2022, 2021), Glosses (Camacho-Collados et al., 2016), Habibi (El-Haj, 2020), HinDialect (Bafna, 2022), HornMT (HornMT, 2023), IITB (Kunchukuttan et al., 2018), IndicNLP (Nakazawa et al., 2021), Indiccorp (Kakwani et al., 2020), isiZulu (Mabuya et al., 2023), JParaCrawl (Morishita et al., 2020), kinyarwandaSMT (Niyongabo, 2023), LeipzigData (Goldhahn et al., 2012), LINDAT (LINDAT, 2023), Lingala Song Lyrics (Lingala Songs, 2023), LyricsTranslate (LyricsTranslate, 2023), mC4 (Raffel et al., 2020), MTData (Gowda et al., 2021), MaCoCu (Bañón et al., 2022), Makerere MT Corpus (Mukiibi et al., 2022), Masakhane Community (Masakhane, 2023), Mburisano Covid Corpus (SADiLaR, 2023a), Menyo20K (Adelani et al., 2021a), Minangkabau corpora (Koto and Koto, 2020), MoT (Palen-Michel et al., 2022), NLLB seed (Costa-jussà et al., 2022), Nart Abkhaz text (Nart, 2023), OPUS (Tiedemann, 2012), OSCAR (Ortiz Suárez et al., 2019), ParaCrawl (Bañón et al., 2020), Parallel Corpora for Ethiopian Languages (Teferra Abate et al., 2018), Phontron (Neubig, 2011), QADI (Abdelali et al., 2021), Quechua-IIC (Zevallos et al., 2022), SLI GalWeb.1.0 (Agerri et al., 2018), Shami (Abu Kwaik et al., 2018), Stanford NLP (Stanford, 2023), StatMT (Koehn, 2023), TICO (Anastasopoulos et al., 2020), TIL (Mirzakhalov et al., 2021), Tatoeba (Tiedemann, 2020), TeDDi (Moran et al., 2022), Tilde (Rozis and Skadiņš, 2017), W2C (Majliš, 2011), WAT (Nakazawa et al., 2022), WikiMatrix (Schwenk et al., 2021), Wikipedia (Wikipedia, 2024), Workshop on NER for South and South East Asian Languages (Singh, 2008), and XLSum (Hasan et al., 2021).
  • MADLAD-400 (Kudugunta et al., 2023):
    CommonCrawl (Common Crawl, 2022).
We start with the corpus from Chang et al. (2023). We then merge the dataset per language with Glot500 for languages that have not yet reached our 1GB maximum (after byte premium scaling). Then, we merge the dataset with MADLAD-400 for languages that have still not reached our 1GB maximum. We also add MADLAD-400 for languages with short average line lengths (less than 25.0 tokens), to make use of MADLAD-400’s longer contiguous sequences. To allow comparisons on popular low-resource language evaluations, we exclude FLORES-200 (Costa-jussà et al., 2022) and AmericasNLI (Ebrahimi et al., 2022) from all dataset merging. For each dataset, we exclude languages that contain only Bible data. Because there is likely significant overlap between different dataset sources, we deduplicate repeated sequences of 100 UTF-8 bytes for each language (Lee et al., 2022).
Language codes.
To enable dataset merging per language, several datasets must be converted to ISO 639-3 language codes and ISO 15924 script codes. In some cases, this introduces ambiguity because datasets can be labeled as individual language codes (e.g. quy_latn for Ayacucho Quechua and quz_latn for Cusco Quechua) or as macrolanguage codes (e.g. que_latn for Quechua). In these cases, we compile both a macrolanguage dataset and individual language datasets. Datasets labeled with individual codes contribute both to their individual dataset and their umbrella macrolanguage dataset; datasets labeled with macrolanguage codes contribute only to the macrolanguage dataset. For example, we have individual quy_latn and quz_latn datasets, both of which contribute to a larger que_latn dataset, which also contains datasets labeled only with que_latn. These ambiguities primarily appear for lower-resource languages.
Additionally, we drop several redundant language codes:
  • We drop ory_orya (Odia) in favor of the macrocode ori_orya because ory_orya is the only individual language within ori_orya for which we have any data.
  • For the same reason, we drop npi_deva (Nepali) in favor of the macrocode nep_deva.
  • For the same reason, we drop swh_latn (Swahili) in favor of the macrocode swa_latn.
  • We drop cmn_hans (Mandarin) in favor of the macrocode zho_hans (Chinese) because the zho_hans data is almost entirely in Mandarin. While less specific, zho_hans is commonly used by other datasets. For other Chinese languages, see their individual codes (e.g. yue_hant for Cantonese). We note that the similar code zho_hant (traditional characters) is not primarily Mandarin.
  • We drop hbs_cyrl and hbs_latn (Serbo-Croatian) because we have the individual languages Serbian (srp_cyrl and srp_latn), Croatian (hrv_latn), and Bosnian (bos_cyrl and bos_latn).
  • We drop the deprecated code ajp_arab (Levantine Arabic) in favor of apc_arab.
  • We drop ber_latn (Berber) because it is a collective code for distinct (and often not mutually intelligible) languages. We keep the constituent individual languages.
  • We drop nah_latn (Nahuatl) because it is a collective code for distinct languages. We keep the constituent individual languages.
After merging, we have a dataset of 547GB of text covering 523 language-script combinations (486 unique language codes, 32 unique script codes).
Byte premiums.
As described in §3.1, we then scale our dataset sizes by estimated byte premiums (Arnett et al., 2024). A byte premium $b$ for a language $L$ indicates that content-matched (i.e. parallel) text in $L$ takes $b\times$ as many UTF-8 bytes to encode as English. We use the Byte Premium Tool (Arnett et al., 2024) to compute or estimate the byte premium for all of our languages. Byte premiums are pre-computed in the tool for high-resource languages. For each novel low-resource language $L$, we use the tool (which uses a linear regression) to predict the byte premium for $L$ based on the character entropy for text in $L$ and the script type for $L$ (alphabet, abjad, abugida, or logography), as recommended for low-resource languages in Arnett et al. (2024). Then, we have an estimated byte premium for every language in our dataset. We clip each byte premium to a minimum of 0.70 and a maximum of 5.00; clipping occurs for only three languages (lzh_hant, wuu_hani $\to$ 0.70, mya_mymr $\to$ 5.00). As described in §3.1, all of our training datasets (both for tokenizers and for the models themselves) are sampled based on size in bytes after byte premium scaling. We drop languages with less than 5MB of text after byte premium scaling.
Dataset statistics.
The resulting 350 Goldfish languages cover five continents, 28 top-level language families (Hammarström et al., 2023), and 32 scripts (writing systems). All languages for which Goldfish models are available are listed in Table LABEL:tab:lang-list. We include the language name, ISO 639-3 language code, ISO 15924 script code, estimated byte premium, dataset size after byte premium scaling, dataset size in tokens, and proportion of the dataset from each of our four largest sources. Raw dataset sizes before byte premium scaling can be obtained by multiplying the dataset size after byte premium scaling by the estimated byte premium. Source dataset proportions are reported before deduplication. The reported dataset sizes reflect the dataset for the Goldfish model trained on the maximum amount of data for that language (the 1GB-dataset Goldfish when available, otherwise the full-dataset Goldfish). Reported token counts use the tokenizer for the largest Goldfish model for that language. All dataset statistics can be downloaded at https://github.com/tylerachang/goldfish.
A.2 Pre-Training Details
As described in §3.2, we train monolingual language models for five dataset sizes when available after byte premium scaling: 5MB, 10MB, 100MB, 1GB, and full. The full dataset size (including all available data) is only included if a 1GB dataset is not available for a language. In total, the Goldfish include 350 5MB-dataset models, 288 10MB-dataset models, 166 100MB-dataset models, 83 1GB-dataset models, and 267 full-dataset models (1154 models total). Full hyperparameters are reported in Table 4.
Hyperparameter 5MB,10MB 100MB,1GB,full
Total parameters 39M 125M
Layers 4 12
Embedding size 512 768
Hidden size 512 768
Intermediate hidden size 2048 3072
Attention heads 8 12
Attention head size 64 64
Learning rate 1e-4
Batch size 5MB: 4, 10MB: 8,
100MB: 32, 1GB: 64
Epochs 10
Activation function GELU
Max sequence length 512
Position embedding Absolute
Learning rate decay Linear
Warmup steps 10% of pre-training
Adam $\epsilon$ 1e-6
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.999
Dropout 0.1
Attention dropout 0.1
Table 4: Pre-training hyperparameters for Goldfish trained on different dataset sizes (Devlin et al., 2019; Turc et al., 2019; Radford et al., 2018).
Tokenizers.
All tokenizers are trained with vocabulary size 50K (Liu et al., 2019) on the same dataset size as their corresponding model (including byte premium scaling). We use SentencePiece tokenizers (Kudo and Richardson, 2018) with training text randomly sampled from the dataset for the desired language. To avoid memory errors, we limit tokenizer training text to 100MB after byte premium scaling. After tokenizer training, we tokenize each training dataset, concatenating text lines such that each sequence contains exactly 512 tokens. We run tokenization before shuffling and sampling to the desired dataset sizes, so our sequences of 512 tokens preserve contiguous text where possible, although several of our source corpora only exist in shuffled form. Finally, we sample our tokenized datasets to 5MB, 10MB, 100MB, and 1GB after byte premium scaling. [^6] When de-tokenized, the tokenized datasets result in slightly smaller datasets than the original text datasets, because the tokenizer truncates lines to create 512-token sequences. All reported dataset sizes account for this truncation.
Architectures.
All of our models use the GPT-2 architecture (Radford et al., 2019), changing only the number of layers, attention heads, and embedding sizes as in Turc et al. (2019). For the 100MB-, 1GB-, and full-dataset models, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018) (similar to BERT-base and RoBERTa; Devlin et al., 2019; Liu et al., 2019). Because smaller models perform similarly to larger models in low-resource scenarios (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes.
Training hyperparameters.
Language models are pre-trained using the Hugging Face Transformers library (Wolf et al., 2020) and code from Chang and Bergen (2022). We refrain from extensive hyperparameter tuning to avoid biasing our hyperparameters towards English (or any other selected tuning language). Instead, we adopt hyperparameters from previous work with minimal modifications. To match the setup of our models and to prevent overfitting, we select hyperparameters based on models with fairly small training datasets relative to modern standards. Specifically, following BERT (Devlin et al., 2019), we use learning rate 1e-4 for the 125M-parameter models (the same as RoBERTa for small batch sizes; Liu et al., 2019; GPT-1 uses learning rate 2.5e-4; Radford et al., 2018). Based on initial results using randomly-sampled languages, we find that learning rate 1e-4 also works well for the 39M-parameter models; this is in line with Chang et al. (2023), who find that learning rate 2e-4 works well for small models, and smaller learning rates reduce the speed of any potential overfitting.
We train each model for 10 epochs of the training data; multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but pre-training on more than 10 epochs often leads to overfitting (increases in eval loss) in the 5MB scenarios. For batch sizes, following GPT-1 (most similar to our models; Radford et al., 2018), we use batch size 64 (64$\times$512 = 32K tokens) for the 1GB-dataset models. We find that these larger batch sizes lead to overfitting for small datasets, so we use batch sizes 4, 8, and 32 for 5MB-, 10MB-, and 100MB-dataset models respectively (determined based on initial experiments with randomly-sampled languages). These correspond to batches of 2K, 4K, or 16K tokens. For full-dataset models, we use the batch size that would be used if rounding the dataset size down to 5MB, 10MB, or 100MB (recall that we do not train a full-dataset model when the 1GB dataset is available for a language).
Compute costs.
All language model pre-training runs together take a total of $1.65\times 10^{20}$ FLOPs. This is less than 1/1900×1/1900$\times$1 / 1900 × the computation used to train the original 175B-parameter GPT-3 model (Brown et al., 2020; $3.14\times 10^{23}$ FLOPs). Models are each trained on one NVIDIA GeForce GTX TITAN X, GeForce RTX 2080 Ti, TITAN Xp, Quadro P6000, RTX A4500, RTX A5000, or RTX A6000 GPU. In total, Goldfish pre-training takes the equivalent of approximately 15600 A6000 GPU hours. Inference for FLORES perplexities and reasoning benchmarks takes approximately 250 A6000 GPU hours (primarily due to the large multilingual models used for comparison). Dataset merging, deduplication, and tokenization takes approximately 1600 CPU core hours.
A.3 FLORES Evaluation Details
[…]
FLORES log-perplexities for all models and languages are reported in Table LABEL:tab:ppl-table. For Goldfish models, we report the log-perplexity for the model trained on the largest dataset for the language (i.e. the 1GB-dataset model when available, otherwise the full-dataset model). Log-perplexities of the 5MB-, 10MB-, 100MB-, and 1GB-dataset models specifically are available at https://github.com/tylerachang/goldfish.
Bigram model details.
For each FLORES language, we train a bigram model on the entire Goldfish dataset for that language, up to 1GB after byte premium scaling §3.1. The bigram model computes the probability of each token $w_{i}$ as $P(w_{i}|w_{i-1})$, computed based on raw bigram counts in the tokenized Goldfish dataset. The tokenizer is the same as the Goldfish tokenizer for that dataset (i.e. the 1GB-dataset model when available, or the full-dataset model). When a bigram is not observed in the dataset, we use backoff to unigram probability with a penalty multiplier of $\lambda=0.40$ (i.e. “stupid backoff”; Brants et al., 2007). We do not consider $n$-grams for $n>2$ because those $n$-grams often resort to backoff and are therefore much more sensitive to the backoff penalty term $\lambda$.
Ambiguous or missing languages.
Several of the FLORES and Belebele languages are either missing from Goldfish or have multiple possible Goldfish available (e.g. either the macrolanguage que_latn or individual language quy_latn for FLORES language quy_latn). We make the following substitutions:
  • taq_tfng $\to$ None,
    tzm_tfng $\to$ None.
    None of the language models evaluated are trained on these languages, and no Goldfish are trained with the Tifinagh (tfng) script.
  • awa_deva $\to$ hin_deva,
    kam_latn $\to$ kik_latn.
    kas_arab $\to$ urd_arab,
    mni_beng $\to$ ben_beng,
    nus_latn $\to$ din_latn,
    taq_latn $\to$ kab_latn,
    Here, we use the closest relative in Goldfish that uses the same script.
  • ace_arab $\to$ urd_arab,
    arb_latn $\to$ mlt_latn,
    ben_latn $\to$ hin_latn,
    bjn_arab $\to$ urd_arab,
    min_arab $\to$ urd_arab,
    npi_latn $\to$ hin_latn,
    sin_latn $\to$ hin_latn,
    urd_latn $\to$ hin_latn,
    These are languages that are missing from Goldfish and that are written in a nonstandard script for the language (e.g. Arabic in Latin script). We use the closest relative in Goldfish that uses that script.
  • acm_arab $\to$ arb_arab,
    acq_arab $\to$ arb_arab,
    aeb_arab $\to$ arb_arab,
    ajp_arab $\to$ arb_arab,
    als_latn $\to$ sqi_latn,
    ars_arab $\to$ arb_arab,
    ary_arab $\to$ arb_arab,
    ayr_latn $\to$ aym_latn,
    azb_arab $\to$ aze_arab,
    azj_latn $\to$ aze_latn,
    dik_latn $\to$ din_latn,
    gaz_latn $\to$ orm_latn,
    khk_cyrl $\to$ mon_cyrl,
    kmr_latn $\to$ kur_latn,
    lvs_latn $\to$ lav_latn,
    npi_deva $\to$ nep_deva,
    ory_orya $\to$ ori_orya,
    pbt_arab $\to$ pus_arab,
    plt_latn $\to$ mlg_latn,
    quy_latn $\to$ que_latn,
    swh_latn $\to$ swa_latn,
    uzn_latn $\to$ uzb_latn,
    ydd_hebr $\to$ yid_hebr,
    yue_hant $\to$ zho_hant,
    zsm_latn $\to$ msa_latn,
    These languages map to multiple different Goldfish languages or are individual languages within a macrolanguage code included in Goldfish. When the option is available, we use the Goldfish language with more data.
A.4 Reasoning Task Details
In §5, we evaluate the Goldfish models, XGLM 4.5B, XGLM 7.5B, BLOOM 7.1B, and MaLA-500 10B on:
  • Non-English Belebele (121 languages, reading comprehension; Bandarkar et al., 2024). For languages that are ambiguous or missing from Goldfish, we use the same language code mapping as in §A.3. Each Belebele example consists of a passage, a question, and four candidate answers. We evaluate model accuracy in selecting the correct answer by computing text probabilities for each “[passage] [question] [answer_option]”. No model exceeds 41% accuracy for any language (random chance 25%).
  • XCOPA (11 languages, commonsense reasoning; Ponti et al., 2020). Each example consists of a premise sentence and two possible causes or effects (i.e. answer options). We use the task format and evaluation implementation in Gao et al. (2023). This selects answers based on a model’s computed text probabilities for each “[premise] [connecting_word] [cause/effect_option]”, where the connecting word is the translation of “because” (for causes) or “therefore” (for effects).
  • Non-English XStoryCloze (10 languages, story commonsense; Lin et al., 2022). Each example consists of a context story and two possible story completions. We use the task format and evaluation implementation in Gao et al. (2023). This selects answers based on a model’s computed text probabilities for each “[story_context] [completion_option]”.
All models are evaluated zero shot with no fine-tuning. Results per language are available at https://github.com/tylerachang/goldfish.
Table 5: FLORES log-perplexity score ($\downarrow$) for each model and FLORES language. Parentheses indicate that the model is not trained specifically on that language.
Language   FLORES log-perplexity
Goldfish Bigram XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
ace_arab (287.55) (365.59) (260.50) (263.32) (232.76) (251.92)
ace_latn 144.62 169.33 (202.67) (208.64) (198.50) 133.10
acm_arab (96.60) (124.48) (93.25) (88.91) (85.28) 102.65
acq_arab (95.54) (124.94) (92.14) (88.09) (82.88) (105.38)
aeb_arab (116.01) (140.30) (113.84) (108.98) (102.76) (120.88)
afr_latn 79.88 115.57 79.54 (162.73) (153.05) 85.61
ajp_arab (98.92) (125.28) (96.25) (91.49) (85.73) 103.56
aka_latn 132.48 162.51 (234.93) (239.68) 187.66 128.37
als_latn 77.28 119.91 76.37 (220.78) (178.58) 89.49
amh_ethi 83.36 111.87 110.54 (266.96) (195.14) 108.99
apc_arab 173.18 179.37 (99.97) (94.47) (87.38) 106.58
arb_arab 82.43 117.61 79.33 75.07 69.24 96.84
arb_latn (245.97) (346.80) 211.13 (226.14) (221.60) (229.07)
ars_arab (83.68) (118.50) (80.72) (76.51) (70.75) (98.25)
ary_arab (128.66) (155.55) (131.80) (125.88) (114.30) 123.30
arz_arab 116.98 146.50 (98.72) (92.96) (87.42) 112.84
asm_beng 93.78 118.79 135.60 (227.40) 113.74 108.20
ast_latn 86.86 118.36 (113.30) (112.31) (99.39) 82.54
awa_deva (128.70) (169.05) (148.81) (135.84) (123.28) (141.50)
ayr_latn 123.33 148.81 (239.30) (231.82) (228.59) 146.85
azb_arab 154.24 185.83 163.32 (218.12) (225.63) 165.85
azj_latn 74.28 106.31 75.33 (183.10) (185.84) 84.13
bak_cyrl 79.24 108.34 (259.93) (272.19) (209.63) 92.72
bam_latn 158.88 175.25 188.31 (212.45) 203.82 143.14
ban_latn 121.16 137.57 (154.09) (183.02) (188.39) 114.41
bel_cyrl 79.50 124.53 81.13 (226.99) (211.85) 91.22
bem_latn 150.39 174.20 (218.51) (188.14) (237.98) 158.61
ben_beng 80.71 109.08 90.55 79.78 76.54 94.62
bho_deva 121.88 138.45 (168.84) (149.88) (120.97) (130.48)
bjn_arab (297.57) (383.82) (260.05) (261.94) (238.42) (257.15)
bjn_latn 111.90 140.40 (151.76) (150.82) (153.76) 104.37
bod_tibt 118.58 142.65 (134.47) (137.26) (206.06) 120.94
bos_latn 73.84 113.42 63.67 (155.27) (135.77) 67.75
bug_latn 172.63 181.80 (206.57) (214.07) (218.78) (211.21)
bul_cyrl 71.36 109.94 64.51 59.03 (148.72) 70.69
cat_latn 76.30 115.98 65.93 60.65 60.55 66.57
ceb_latn 94.04 125.05 (111.22) (192.82) (171.75) 103.37
ces_latn 75.07 115.11 63.68 (154.38) (140.67) 71.35
cjk_latn 219.90 239.42 (212.05) (211.78) (220.85) 189.97
ckb_arab 89.11 121.96 (112.62) (290.64) (199.22) 107.81
crh_latn 105.56 127.10 (175.84) (185.40) (180.14) 119.20
cym_latn 77.50 114.42 97.27 (226.16) (204.15) 100.22
dan_latn 74.09 111.50 60.26 (122.87) (129.58) 67.08
deu_latn 73.91 118.06 57.93 57.08 (98.55) 62.85
dik_latn 152.14 169.32 (197.55) (202.82) (216.99) (225.78)
dyu_latn 183.05 210.69 (209.72) (216.63) (209.24) 189.79
dzo_tibt 125.44 144.63 (213.72) (217.28) (238.66) 110.58
ell_grek 78.38 119.29 68.55 65.99 (157.23) 93.41
eng_latn 68.73 103.16 51.10 50.39 50.56 48.43
epo_latn 75.57 110.48 118.68 (175.58) (149.97) 88.92
est_latn 73.18 109.48 71.72 60.94 (173.50) 96.32
eus_latn 70.76 103.94 112.97 70.24 70.78 101.32
ewe_latn 128.71 144.86 (181.68) (202.81) (228.49) 137.60
fao_latn 89.22 120.77 (182.64) (216.89) (180.74) 100.02
fij_latn 107.24 126.01 (186.08) (136.05) (215.93) (121.46)
fin_latn 75.34 114.20 63.80 55.50 (164.25) 82.10
fon_latn 190.75 205.52 (209.87) (255.39) 239.02 190.66
fra_latn 70.55 116.07 56.67 55.76 53.43 59.64
fur_latn 114.09 131.26 (203.28) (198.54) (185.86) 107.32
fuv_latn 165.36 188.23 (188.00) (201.33) (192.94) (184.41)
gaz_latn 120.21 202.78 (163.96) (260.02) (262.77) (144.71)
gla_latn 97.03 133.97 172.79 (245.08) (209.10) 120.94
gle_latn 79.48 118.25 150.60 (228.43) (204.19) 106.38
glg_latn 76.34 114.53 86.26 (107.84) (87.42) 74.13
grn_latn 121.55 141.64 193.46 (230.52) (225.65) 119.96
guj_gujr 84.11 110.95 100.35 (276.27) 96.00 97.04
hat_latn 89.82 114.98 114.15 85.51 (163.36) 102.63
hau_latn 86.28 116.50 115.81 (215.68) (215.07) 107.27
heb_hebr 77.17 108.77 73.71 (175.48) (147.78) 111.24
hin_deva 76.64 110.74 79.74 71.13 70.06 88.97
hne_deva 141.43 149.99 (164.54) (153.61) (144.13) 110.53
hrv_latn 71.16 110.07 61.93 (153.35) (135.73) 66.55
hun_latn 75.05 113.40 64.13 (176.50) (164.44) 78.54
hye_armn 78.37 115.07 85.23 (259.42) (224.51) 94.83
ibo_latn 110.84 135.25 147.33 (248.34) 149.46 123.89
ilo_latn 111.06 129.57 (133.69) (227.74) (209.68) 115.75
ind_latn 72.11 101.11 62.46 60.13 59.44 65.23
isl_latn 75.36 113.10 82.56 (200.00) (179.47) 93.17
ita_latn 75.27 116.41 60.06 59.00 (86.91) 62.98
jav_latn 90.19 112.98 102.58 (161.56) (154.48) 102.21
jpn_jpan 68.51 100.99 63.11 61.53 (93.28) 68.24
kab_latn 134.05 154.27 (214.03) (256.69) (215.19) 159.99
kac_latn 128.22 137.82 (187.08) (248.17) (239.55) 146.59
kam_latn (240.45) (264.47) (186.89) (202.02) (200.47) 172.19
kan_knda 76.34 103.10 92.36 (268.95) 93.76 97.03
kas_arab (252.81) (304.08) (276.83) (267.41) (245.94) (260.94)
kas_deva 221.04 240.29 (235.98) (228.86) (231.57) (246.51)
kat_geor 72.44 108.39 82.24 (261.79) (236.64) 96.53
kaz_cyrl 73.52 102.69 76.14 (192.47) (186.61) 89.71
kbp_latn 145.05 159.84 (217.85) (264.27) (286.31) 143.65
kea_latn 145.29 157.19 (180.88) (187.08) (182.18) 135.49
khk_cyrl 77.40 103.60 87.49 (238.21) (192.81) 97.14
khm_khmr 98.82 139.14 114.88 (323.85) (240.85) 117.47
kik_latn 165.85 177.82 (222.18) (244.09) 227.12 148.48
kin_latn 87.00 118.79 (227.82) (225.85) 135.94 113.52
kir_cyrl 70.91 99.21 114.55 (224.69) (202.97) 94.96
kmb_latn 179.18 199.36 (205.24) (196.76) (216.45) 168.66
kmr_latn 99.93 130.74 155.62 (234.52) (213.35) 120.91
knc_arab 181.38 274.16 (223.08) (222.78) (214.43) (228.52)
knc_latn 170.17 206.34 (229.82) (227.51) (239.24) (242.18)
kon_latn 132.91 143.02 182.94 (190.80) (189.74) (126.76)
kor_hang 72.23 102.86 69.83 63.54 (122.61) 73.28
lao_laoo 91.00 120.61 110.10 (268.95) (231.49) 107.59
lij_latn 140.38 168.14 (198.93) (195.96) (193.23) 122.99
lim_latn 123.42 154.99 (180.85) (201.88) (192.90) (116.94)
lin_latn 106.45 122.73 144.57 (183.46) 167.57 116.17
lit_latn 71.55 110.06 67.23 (188.51) (163.54) 92.94
lmo_latn 162.18 201.95 (203.26) (199.20) (198.36) 152.10
ltg_latn 121.88 138.23 (217.38) (241.02) (229.77) (227.25)
ltz_latn 85.00 123.40 (154.55) (149.83) (188.39) 103.80
lua_latn 152.56 162.72 (175.04) (167.40) (201.83) 147.08
lug_latn 118.73 139.82 168.52 (213.70) 165.48 141.39
luo_latn 139.03 155.32 (210.18) (218.56) (222.85) 163.70
lus_latn 95.13 126.88 (157.04) (217.70) (211.72) 131.48
lvs_latn 70.94 110.34 70.45 (187.39) (179.57) 86.42
mag_deva 126.54 139.42 (156.88) (143.63) (128.48) (133.86)
mai_deva 123.50 142.56 (174.69) (165.19) (126.39) 102.51
mal_mlym 80.56 111.47 92.78 (221.87) 88.15 99.60
mar_deva 82.32 110.97 94.21 (200.05) 92.70 100.33
min_arab (308.31) (399.00) (269.74) (273.93) (254.70) (275.42)
min_latn 108.38 128.37 (167.37) (164.01) (160.32) 125.86
mkd_cyrl 73.08 109.43 73.22 (134.89) (162.92) 76.94
mlt_latn 83.60 125.90 (280.75) (279.00) (237.59) 90.70
mni_beng (176.58) (274.13) (279.36) (271.00) (188.18) (275.78)
mos_latn 187.64 198.01 (228.35) (236.65) (241.18) 188.11
mri_latn 97.39 130.22 (191.98) (187.99) (180.89) 109.67
mya_mymr 86.45 125.64 119.52 90.49 (224.44) 121.00
nld_latn 71.72 112.22 60.19 (111.54) (115.43) 65.13
nno_latn 80.80 114.77 (73.23) (153.12) (151.17) 76.16
nob_latn 76.13 109.96 64.70 (122.09) (131.71) 68.04
npi_deva 82.42 111.11 90.72 (193.40) 86.21 86.53
nso_latn 119.03 139.17 (166.19) (234.76) 181.13 123.81
nus_latn (217.98) (276.12) (259.51) (266.18) (293.48) (319.71)
nya_latn 97.55 121.06 (213.77) (202.58) 180.39 118.77
oci_latn 97.75 135.49 (152.88) (126.52) (129.14) 93.79
ory_orya 87.64 114.38 (113.19) (387.80) (103.65) 105.48
pag_latn 122.99 139.15 (175.08) (184.33) (184.39) 120.19
pan_guru 85.22 116.57 107.69 (298.41) 99.43 101.39
pap_latn 94.63 126.00 (166.14) (177.79) (183.93) 118.24
pbt_arab 104.87 136.61 120.56 (263.66) (210.05) 130.62
pes_arab 75.11 112.64 70.67 (152.72) (145.67) 84.23
plt_latn 87.89 121.62 116.37 (236.48) (176.35) 102.04
pol_latn 72.87 114.38 61.18 (146.68) (130.92) 69.19
por_latn 72.38 110.34 59.80 57.79 55.73 63.06
prs_arab 96.31 120.44 (80.74) (150.59) (141.22) 83.70
quy_latn 121.48 144.34 185.50 125.42 (196.23) 132.70
ron_latn 75.68 118.45 62.78 (151.24) (135.97) 70.39
run_latn 112.80 135.64 (229.50) (226.86) 152.96 127.91
rus_cyrl 73.59 117.13 58.22 57.38 (110.95) 65.18
sag_latn 162.70 167.88 (182.49) (171.77) (196.64) 150.49
san_deva 134.36 156.91 167.12 (188.71) (167.33) 140.17
sat_olck 148.03 156.26 (217.91) (217.35) (298.99) 124.11
scn_latn 124.37 157.13 (172.60) (187.74) (175.53) 108.90
shn_mymr 162.52 182.33 (253.11) (475.53) (373.38) (341.88)
sin_sinh 83.19 114.20 97.91 (304.98) (235.47) 106.82
slk_latn 72.22 113.10 63.60 (181.84) (157.25) 78.82
slv_latn 71.40 111.05 64.99 (176.07) (147.81) 75.62
smo_latn 103.20 142.22 (197.35) (213.38) (214.55) 120.15
sna_latn 93.53 123.08 (228.56) (225.74) 180.82 125.19
snd_arab 91.04 116.69 150.37 (257.35) (212.74) 117.36
som_latn 104.72 138.92 125.40 (245.50) (229.18) 135.24
sot_latn 99.85 142.34 (171.86) (235.46) 192.76 122.30
spa_latn 77.45 116.81 63.10 61.91 59.82 66.71
srd_latn 116.75 135.26 (202.61) (200.61) (191.63) 106.39
srp_cyrl 76.66 116.32 72.33 (175.02) (149.07) 71.86
ssw_latn 124.51 143.31 186.61 (232.90) (220.26) 133.78
sun_latn 91.07 116.80 109.40 (167.76) (162.37) 98.90
swe_latn 73.76 109.95 62.27 (106.34) (126.73) 65.90
swh_latn 79.45 109.31 89.60 76.81 98.22 95.63
szl_latn 131.79 157.70 (184.61) (229.41) (208.14) 122.68
tam_taml 79.87 107.64 88.40 78.39 79.09 95.70
taq_latn (241.82) (267.82) (216.97) (231.56) (222.07) (225.98)
taq_tfng None None (301.75) (289.98) (261.59) (396.24)
tat_cyrl 76.67 107.20 (222.35) (227.31) (192.83) 88.82
tel_telu 80.76 106.79 89.99 77.70 88.52 95.52
tgk_cyrl 79.61 113.99 (267.56) (281.19) (213.45) 100.89
tgl_latn 86.11 123.04 91.13 (168.86) (162.76) 89.23
tha_thai 73.84 102.82 71.03 64.31 (177.26) 87.04
tir_ethi 107.10 134.77 142.22 (300.08) (258.62) 133.08
tpi_latn 141.91 164.61 (218.27) (212.39) (197.37) 109.12
tsn_latn 111.75 143.35 184.89 (240.07) 186.95 127.47
tso_latn 111.93 133.30 (236.13) (239.86) 205.36 130.50
tuk_latn 81.23 107.98 (239.28) (256.07) (198.90) 114.93
tum_latn 132.63 154.17 (233.00) (187.16) 237.42 141.81
tur_latn 69.75 100.52 64.44 61.13 (130.11) 86.26
twi_latn 131.51 148.42 (211.42) (216.89) 174.63 128.30
tzm_tfng None None (206.77) (206.42) (243.59) (332.15)
uig_arab 75.68 105.62 (317.09) (367.46) (206.46) 105.21
ukr_cyrl 76.60 116.07 64.77 (129.73) (145.72) 72.19
umb_latn 182.34 211.34 (199.72) (209.91) (221.59) 174.09
urd_arab 83.96 117.74 90.05 79.04 85.58 98.80
uzn_latn 71.09 107.26 112.09 (243.88) (217.18) 96.67
vec_latn 114.88 147.99 (161.71) (155.44) (160.51) 108.87
vie_latn 77.66 120.08 68.06 64.89 61.00 74.36
war_latn 118.02 153.85 (161.42) (203.73) (174.18) 132.17
wol_latn 141.12 158.76 202.72 (225.23) 167.95 161.96
xho_latn 93.68 121.39 144.06 (216.55) 155.76 122.42
ydd_hebr 109.90 144.63 (260.53) (286.49) (210.28) 128.80
yor_latn 123.24 167.30 174.33 (246.05) 154.45 148.67
yue_hant 90.03 121.49 (70.31) (86.41) 61.81 69.76
zho_hans 78.92 121.82 66.34 65.42 59.08 70.09
zho_hant 93.56 125.41 (72.53) (89.36) 63.06 (75.40)
zsm_latn 73.25 100.45 67.03 (83.84) (76.22) 72.90
zul_latn 87.90 118.40 135.47 (218.33) 186.47 115.47
Table 6: Goldfish languages with corresponding dataset sizes.
Language   Language Script Byte Scaled Tokens Dataset Proportions
(ISO 639-3) (ISO 15924) Premium MB
  OSCAR
  NLLB
  MADLAD-400
  Glot500
  Other
Afrikaans afr latn 1.04 1000.00 239682048
Amharic amh ethi 1.72 1000.00 211767808
Standard Arabic arb arab 1.47 1000.00 196197376
Azerbaijani aze latn 1.30 1000.00 233091584
Belarusian bel cyrl 2.01 1000.00 254138368
Bengali ben beng 2.43 1000.00 194737152
Bosnian bos cyrl 1.15 1000.00 232501760
Bosnian bos latn 0.97 1000.00 228266496
Bulgarian bul cyrl 1.81 1000.00 224346112
Catalan cat latn 1.09 1000.00 238915072
Czech ces latn 1.04 1000.00 206113280
Welsh cym latn 1.03 1000.00 236230144
Danish dan latn 1.02 1000.00 208085504
German deu latn 1.05 1000.00 210817024
Modern Greek ell grek 1.97 1000.00 238704128
English eng latn 1.00 1000.00 213977088
Esperanto epo latn 1.00 1000.00 231384576
Estonian est latn 0.97 1000.00 189518336
Basque eus latn 1.06 1000.00 209921536
Persian fas arab 1.59 1000.00 244359680
Filipino fil latn 1.33 1000.00 274955776
Finnish fin latn 1.06 1000.00 186050560
French fra latn 1.17 1000.00 251415552
Galician glg latn 1.06 1000.00 222080000
Gujarati guj gujr 2.16 1000.00 193794560
Hausa hau latn 1.18 1000.00 277416448
Hebrew heb hebr 1.36 1000.00 192904704
Hindi hin deva 2.37 1000.00 228020736
Croatian hrv latn 0.99 1000.00 219422208
Hungarian hun latn 1.02 1000.00 191089664
Armenian hye armn 1.72 1000.00 203630592
Indonesian ind latn 1.18 1000.00 210432000
Icelandic isl latn 1.15 1000.00 236872704
Italian ita latn 1.07 1000.00 216099840
Japanese jpn jpan 1.32 1000.00 219063296
Kara-Kalpak kaa cyrl 1.92 1000.00 212100608
Kannada kan knda 2.64 1000.00 212683264
Georgian kat geor 4.34 1000.00 354762752
Kazakh kaz cyrl 1.76 1000.00 199970304
Kirghiz kir cyrl 1.96 1000.00 223066112
Korean kor hang 1.29 1000.00 227021824
Latin lat latn 0.88 1000.00 188774912
Latvian lav latn 1.29 1000.00 243401728
Lithuanian lit latn 1.03 1000.00 201228800
Malayalam mal mlym 2.88 1000.00 244708864
Marathi mar deva 2.48 1000.00 206630400
Macedonian mkd cyrl 1.83 1000.00 221346304
Maltese mlt latn 1.09 1000.00 283158528
Mongolian mon cyrl 1.78 1000.00 205737472
Malay msa latn 1.29 1000.00 236371456
Nepali nep deva 2.63 1000.00 215368192
Dutch nld latn 1.05 1000.00 216978432
Norwegian Bokmål nob latn 1.00 1000.00 205949952
Norwegian nor latn 1.13 1000.00 255482880
Panjabi pan guru 2.22 1000.00 215775232
Iranian Persian pes arab 1.60 1000.00 215946240
Polish pol latn 1.08 1000.00 216235008
Portuguese por latn 1.10 1000.00 225242112
Pushto pus arab 1.59 1000.00 237871616
Romanian ron latn 1.12 1000.00 230580224
Russian rus cyrl 1.82 1000.00 220467712
Sinhala sin sinh 2.45 1000.00 233098752
Slovak slk latn 1.04 1000.00 211206144
Slovenian slv latn 0.97 1000.00 198052864
Somali som latn 1.42 1000.00 302652928
Spanish spa latn 1.08 1000.00 221790720
Albanian sqi latn 1.34 1000.00 274664448
Serbian srp cyrl 1.42 1000.00 184423424
Serbian srp latn 0.83 1000.00 207482368
Swahili swa latn 1.26 1000.00 260033024
Swedish swe latn 1.02 1000.00 206359552
Tamil tam taml 2.73 1000.00 200523264
Tatar tat cyrl 1.85 1000.00 232933888
Telugu tel telu 2.62 1000.00 209365504
Tajik tgk cyrl 1.75 1000.00 216990208
Tagalog tgl latn 1.12 1000.00 245370880
Thai tha thai 2.74 1000.00 205872640
Turkish tur latn 1.04 1000.00 186848768
Ukrainian ukr cyrl 1.75 1000.00 215392768
Urdu urd arab 1.71 1000.00 247899648
Uzbek uzb latn 1.23 1000.00 261058560
Vietnamese vie latn 1.35 1000.00 262306304
Chinese zho hans 0.94 1000.00 206204416
Irish gle latn 1.98 976.70 404823040
Kurdish kur arab 1.57 902.39 196483584
Standard Malay zsm latn 1.14 859.52 185929728
Central Kurdish ckb arab 1.65 838.87 190565888
Kinyarwanda kin latn 1.13 810.96 193561088
Haitian hat latn 0.97 775.80 185333248
Odia ori orya 2.60 774.55 165528576
Zulu zul latn 1.16 764.14 199965696
Burmese mya mymr 5.00 762.14 315374592
Central Khmer khm khmr 3.90 742.37 235559424
Malagasy mlg latn 1.27 720.80 210497024
Kurdish kur latn 1.29 685.53 189872128
Dhivehi div thaa 2.00 634.02 114510336
Shona sna latn 1.12 608.11 151712256
Luxembourgish ltz latn 1.23 579.07 160200192
Sundanese sun latn 1.10 577.96 142266368
Scottish Gaelic gla latn 0.99 558.84 123736064
Cebuano ceb latn 1.11 540.21 140301312
Lao lao laoo 2.71 532.98 124077056
Uzbek uzb cyrl 1.98 525.51 110868992
Yoruba yor latn 1.37 502.55 155829248
Norwegian Nynorsk nno latn 1.03 498.93 116016128
Xhosa xho latn 1.20 477.36 127885824
Western Frisian fry latn 1.23 472.81 133072384
Javanese jav latn 1.15 465.58 115332096
Sindhi snd arab 1.59 459.14 114626048
Maori mri latn 1.18 450.17 136011776
Yiddish yid hebr 1.55 446.04 85695488
Nyanja nya latn 1.21 444.13 112440832
Corsican cos latn 1.18 414.00 126150656
Faroese fao latn 1.16 400.34 96587776
Bashkir bak cyrl 2.27 398.36 118369280
Uighur uig arab 2.31 397.21 104039936
Igbo ibo latn 1.35 388.31 119706112
Modern Greek ell latn 1.24 376.42 92225536
Occitan oci latn 1.01 375.38 99783680
Plateau Malagasy plt latn 1.15 370.58 97517568
Assamese asm beng 2.53 348.88 77216256
Hmong hmn latn 1.19 345.97 100051968
Tosk Albanian als latn 1.17 336.30 87609344
Southern Sotho sot latn 1.17 332.91 94144000
Samoan smo latn 1.18 314.93 101910016
Azerbaijani aze arab 1.20 267.26 56526848
Hawaiian haw latn 1.11 260.95 86747136
Chuvash chv cyrl 1.80 256.36 84293120
Papiamento pap latn 1.00 255.51 60037632
Tigrinya tir ethi 1.76 252.98 56515072
Asturian ast latn 1.75 225.68 93333504
Southern Pashto pbt arab 1.74 225.11 60608000
Central Kanuri knc arab 2.50 221.65 237422592
Lushai lus latn 1.17 213.03 62735360
Northern Uzbek uzn cyrl 2.01 208.92 44960768
Yakut sah cyrl 1.88 206.06 47289344
Ancient Greek grc grek 1.77 205.45 47620608
Turkmen tuk latn 1.79 186.44 57201664
Chinese zho hant 0.99 177.32 42692096
Waray war latn 1.09 175.25 48998912
Kara-Kalpak kaa latn 1.23 165.22 38767104
Breton bre latn 1.01 163.11 43437056
Dari prs arab 1.66 162.70 37549568
Venetian vec latn 1.00 150.70 40523776
North Azerbaijani azj latn 1.08 149.82 27041792
Northern Uzbek uzn latn 1.65 145.59 52049408
Limburgan lim latn 1.00 142.31 39700480
Kalaallisut kal latn 1.34 140.44 30082048
Quechua que latn 1.21 139.38 40595968
Oromo orm latn 1.26 137.90 39742976
Ganda lug latn 1.22 132.42 37459968
Tibetan bod tibt 2.62 131.94 23463424
Hindi hin latn 1.26 131.86 37683712
Swiss German gsw latn 1.14 128.81 38605824
Ayacucho Quechua quy latn 1.16 123.58 34850816
Lombard lmo latn 0.94 123.24 35603456
Egyptian Arabic arz arab 1.55 122.38 30322176
Western Panjabi pnb arab 1.41 121.58 30110208
Eastern Yiddish ydd hebr 1.81 120.20 28306432
Sanskrit san deva 2.54 119.34 31856128
Sicilian scn latn 1.04 113.80 32010752
Halh Mongolian khk cyrl 1.80 108.25 23605760
South Azerbaijani azb arab 1.49 107.56 26922496
Walloon wln latn 1.22 102.32 29091328
Tswana tsn latn 1.17 101.85 31488512
Gujarati guj latn 1.19 101.60 24635392
Gilaki glk arab 1.68 98.73 25519104
Iloko ilo latn 1.08 97.44 25450496
Tetum tet latn 1.40 96.03 28032512
Banjar bjn latn 1.17 93.17 25012224
Rundi run latn 1.12 90.59 23721984
Romansh roh latn 1.27 86.73 23623680
Chechen che cyrl 1.83 86.11 23590400
West Central Oromo gaz latn 1.33 79.04 25565184
Yue Chinese yue hant 0.86 78.42 16084992
Low German nds latn 1.14 75.35 20312064
Minangkabau min latn 0.95 75.07 17732608
Inuktitut iku cans 2.16 74.41 13798400
Tsonga tso latn 1.21 71.85 21684224
Achinese ace latn 1.24 71.09 21666816
Tuvinian tyv cyrl 1.86 68.39 15576576
Northern Sami sme latn 1.27 66.64 15802880
Ewe ewe latn 1.08 63.27 18470400
Twi twi latn 1.03 62.79 18900480
Standard Estonian ekk latn 0.99 61.41 12375552
Guarani grn latn 0.99 60.38 15366656
Pedi nso latn 1.12 59.40 17516544
Northern Kurdish kmr latn 1.03 53.71 12299264
Udmurt udm cyrl 1.74 51.77 10932736
Akan aka latn 1.57 49.51 22551040
Mari (Russia) chm cyrl 1.76 49.43 11290624
Mongolian mon latn 1.18 49.21 12692480
Lingala lin latn 1.14 47.33 13213184
Crimean Tatar crh latn 1.31 47.20 12994560
Zaza zza latn 1.20 46.78 14813184
Kabyle kab latn 1.03 45.19 14035456
Min Nan Chinese nan latn 1.15 44.38 16624128
Scots sco latn 1.19 42.97 12578304
Aragonese arg latn 1.19 42.82 12469760
Maithili mai deva 2.39 41.73 11159040
Fon fon latn 1.54 40.84 13993984
Buriat bua cyrl 1.70 39.10 8951808
Ossetian oss cyrl 1.85 38.60 14059008
Pampanga pam latn 1.19 38.14 11270656
Dimli diq latn 0.96 37.98 9935872
Wolof wol latn 1.08 37.32 12005888
Tedim Chin ctd latn 1.30 37.10 11405824
Tumbuka tum latn 1.21 36.69 9842688
Pangasinan pag latn 1.04 36.43 10441728
Fijian fij latn 1.21 35.48 10642944
Standard Latvian lvs latn 1.21 35.42 8333312
Bemba bem latn 1.16 35.35 10177024
Kabardian kbd cyrl 1.78 34.89 9802752
Luo luo latn 1.04 34.50 9859072
Hakha Chin cnh latn 1.32 33.20 10364928
Hiligaynon hil latn 1.35 32.12 9034752
Balinese ban latn 1.27 31.84 9161216
Aymara aym latn 1.21 30.74 9201152
Avaric ava cyrl 1.94 30.73 8009728
Central Aymara ayr latn 1.10 28.37 7641088
Fiji Hindi hif latn 1.28 28.00 8768000
Ligurian lij latn 1.14 27.89 8498176
Eastern Mari mhr cyrl 1.81 27.86 6580224
Bavarian bar latn 1.13 27.68 7961600
Silesian szl latn 1.07 27.04 7593472
Russian rus latn 1.18 26.62 7373824
Ido ido latn 1.18 26.18 7369216
Russia Buriat bxr cyrl 1.59 25.38 6060544
Abkhazian abk cyrl 2.01 25.24 6408192
Sardinian srd latn 1.11 24.71 6834176
Nigerian Pidgin pcm latn 0.95 24.62 5281280
Wu Chinese wuu hani 0.70 24.53 4112384
Fulah ful latn 1.26 24.03 7806464
Bhojpuri bho deva 2.52 23.74 6156800
Betawi bew cyrl 1.74 23.52 5288960
Volapük vol latn 1.13 21.39 6030336
Nigerian Fulfulde fuv latn 1.11 21.23 6159872
Karachay-Balkar krc cyrl 1.87 21.02 4627456
Swati ssw latn 1.14 20.97 5566976
Luba-Lulua lua latn 1.19 20.82 6322688
Friulian fur latn 1.07 20.72 5487616
Khasi kha latn 1.30 20.56 6209536
Telugu tel latn 1.28 20.02 5266432
Iban iba latn 1.30 19.98 5278208
Bikol bik latn 1.27 19.26 5440512
Interlingua ina latn 1.24 19.15 5581824
Latgalian ltg latn 1.00 18.70 4046848
Komi kom cyrl 1.61 18.20 4716032
Querétaro Otomi otq latn 1.25 17.48 5702656
Tonga (Tonga Islands) ton latn 1.27 17.46 6237184
Azerbaijani aze cyrl 1.82 17.12 3627008
Dargwa dar cyrl 2.02 16.99 4506624
Erzya myv cyrl 1.77 16.81 3851776
Piemontese pms latn 1.23 16.75 5307904
Tok Pisin tpi latn 1.18 16.61 5102592
Umbundu umb latn 1.17 16.12 4743168
Sango sag latn 1.16 15.87 4929024
Kabuverdianu kea latn 0.78 15.74 3247616
Adyghe ady cyrl 1.81 15.22 4124160
Literary Chinese lzh hant 0.70 15.20 2767872
Gulf Arabic afb arab 1.37 14.25 3247616
Falam Chin cfm latn 1.32 14.09 4315648
Kabiyè kbp latn 1.44 13.93 4698624
Bambara bam latn 1.26 12.84 4511744
Kachin kac latn 1.35 12.74 4453888
Newari new deva 2.56 12.44 2927616
Syriac syr syrc 1.41 12.17 2641408
Chokwe cjk latn 1.17 12.10 3622400
Dyula dyu latn 1.15 11.94 3849216
Betawi bew latn 1.30 11.84 3186176
Venda ven latn 1.30 11.82 3268608
Dinka din latn 1.24 11.69 4125696
Shan shn mymr 2.82 11.66 2238976
Southern Altai alt cyrl 1.86 11.65 2694144
Southwestern Dinka dik latn 1.12 11.61 3753984
Goan Konkani gom deva 1.74 11.50 2219520
Sranan Tongo srn latn 1.06 11.47 3098112
Yucateco yua latn 1.24 11.41 3645440
Kongo kon latn 1.23 11.32 3549184
Kimbundu kmb latn 1.13 11.09 3359744
Kumyk kum cyrl 1.96 11.04 2208768
Buginese bug latn 1.23 10.72 3269632
Goan Konkani gom latn 1.21 10.38 2806784
Mossi mos latn 1.14 10.37 3537920
Upper Sorbian hsb latn 1.12 10.31 2503680
Lak lbe cyrl 2.01 10.24 2470912
North Ndebele nde latn 0.97 10.17 1766912
Central Kanuri knc latn 1.18 10.07 3433472
Ingush inh cyrl 1.70 9.59 2764800
Zapotec zap latn 1.08 9.58 2395136
Central Bikol bcl latn 1.22 9.49 2638336
Lezghian lez cyrl 1.83 9.38 2358784
Kituba mkw cyrl 1.81 9.37 2266112
Cusco Quechua quz latn 1.30 9.32 2070528
Bishnupriya bpy beng 2.33 9.29 2019328
Mam mam latn 1.34 9.27 3580416
Magahi mag deva 2.56 9.08 2488832
Tzotzil tzo latn 1.49 9.02 3463680
Tamil tam latn 1.27 9.00 2260992
Western Mari mrj cyrl 1.51 8.74 1812992
Brunei Bisaya bsb latn 1.31 8.69 2460672
Chhattisgarhi hne deva 2.17 8.61 2106880
Luba-Katanga lub latn 1.30 8.61 2269184
Kaqchikel cak latn 1.82 8.51 4157952
Santali sat olck 2.80 8.49 2224128
Vlaams vls latn 1.21 8.49 2484736
Kikuyu kik latn 1.29 8.36 2418176
Mirandese mwl latn 1.24 8.12 2293760
Isoko iso latn 1.48 8.11 2638336
Uighur uig latn 1.19 7.88 1662976
Dzongkha dzo tibt 3.26 7.70 2019328
Bashkir bak latn 1.19 7.53 1793024
Dombe dov latn 0.99 7.43 1389056
Madurese mad latn 1.29 7.29 2044416
Levantine Arabic apc arab 1.47 7.06 1687040
Pohnpeian pon latn 0.90 7.02 1412608
Kashmiri kas deva 2.53 6.96 1990656
Paite Chin pck latn 1.32 6.94 2163712
Veps vep latn 1.17 6.89 1751552
Boko (Benin) bqc latn 0.98 6.80 1806336
Neapolitan nap latn 1.23 6.73 2123776
Manx glv latn 1.22 6.63 1939968
Nande nnb latn 1.31 6.49 1764352
Batak Toba bbc latn 1.33 6.48 1846784
Malayalam mal latn 1.27 6.38 1556480
Tiv tiv latn 1.31 6.32 2119168
Cornish cor latn 1.22 6.31 1936896
Khakas kjh cyrl 1.93 6.17 1271808
Moksha mdf cyrl 1.71 6.17 1302016
Kalmyk xal cyrl 1.72 6.05 1474048
Guerrero Nahuatl ngu latn 1.44 5.99 1508864
Klingon tlh latn 1.14 5.91 1741312
Crimean Tatar crh cyrl 1.89 5.86 1265664
Makhuwa-Meetto mgh latn 1.11 5.77 1251328
Sanskrit san latn 0.97 5.72 1164800
Northern Frisian frr latn 1.17 5.68 1594368
Eastern Balochi bgp latn 1.29 5.64 1735680
Carpathian Romani rmc latn 1.02 5.61 1241600
Georgian kat latn 1.20 5.57 1422336
Old English ang latn 1.29 5.47 1671168
Kedah Malay meo latn 1.28 5.44 1670656
Mingrelian xmf geor 2.51 5.44 1367040
Tulu tcy knda 2.67 5.29 1210368
Tandroy-Mahafaly Malagasy tdx latn 1.00 5.23 1303552
Komi-Zyrian kpv cyrl 1.67 5.19 1355776
Lingua Franca Nova lfn latn 1.30 5.12 1593344
Ditammari tbz latn 1.33 5.12 1868800
Nzima nzi latn 1.42 5.07 1514496
Rusyn rue cyrl 1.56 5.03 1160704
Eastern Huasteca Nahuatl nhe latn 1.49 5.02 1268224